OpenStudy (anonymous):

when m is subtracted from n, the result is r. which of the following expresses the result when 2m is added to s? a. s+2n-2r b. s+2n+2r c. 2s+2n-2r d. 2s+2n+2r e. s-2n+2r

OpenStudy (anonymous):

We know that n-m = r We need to isolate m so we can multiply it by 2. Try it.

OpenStudy (anonymous):

Here, I'll start you off. n -m = r n -m +m = r +m n = r+m Can you get m on its own?